a computer manufacturer offers a computer with a choice of 3 types of monitors, 6 types of keyboards, and 3 types of hard drives. How many different computers are offered?
Answer by jim_thompson5910(35256)

There are 3*6*3 = 18*3 = 54 different types of computers